# Length independent generalization bounds for deep SSM architectures with stability constraints

### v1

@@ Appendix B Proofs
$=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{2,\infty}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\sqrt{\sup\limits_{1\leq k\leq T}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right\rVert_{2}^{2}}\right]$
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{2,\infty}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\sqrt{\sum\limits_{k=1}^{T}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right\rVert_{2}^{2}}\right]$
$\leq\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{2,\infty}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\ell^{2,2}(\mathbb{R}^{n_{u}})}\right],$
where the second equality is the definition of the norm, the first inequality is the standard inequality for induced matrix norms.
c) Again, we only need to deal with the term
$\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W\mathbf{u}_{i}\right\rVert_{Y}\right],$
but now $\mathbf{u}_{i}\in\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})$ and $W\mathbf{u}_{i}\in\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})$. We have
$\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W\mathbf{u}_{i}\right\rVert_{Y}\right]=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert W\left(\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right)\right\rVert_{Y}\right]$
$=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\sup\limits_{1\leq k\leq T}\left\lVert W\left(\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right)\right\rVert_{\infty}\right]$
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\sup\limits_{1\leq k\leq T}\left\lVert W\right\rVert_{\infty,\infty}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right\rVert_{\infty}\right]$
$=\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right],$
where the inequality is again the standard inequality for induced matrix norms.
@@
We can see that the calculactions hold if the transformations are restricted to the ball $B_{X_{1}}(r)$ for any choice of $X_{1}$ we consider. The radius can grow as
@@
$\left\lVert W\mathbf{u}+\mathbf{b}\right\rVert_{\ell^{q,q}(\mathbb{R}^{n_{v}})}\leq\left\lVert W\mathbf{u}\right\rVert_{\ell^{q,q}(\mathbb{R}^{n_{v}})}+\left\lVert\underline{\mathbf{b}}\right\rVert_{\ell^{q,q}(\mathbb{R}^{n_{v}})}\leq\left\lVert W\right\rVert_{p,q}\left\lVert\mathbf{u}\right\rVert_{\ell^{p,p}(\mathbb{R}^{n_{u}})}+\left\lVert\mathbf{b}\right\rVert_{q}.$
Proof [Proof of Lemma 11] Let $Z\subseteq X_{1}^{N}$ and $\tilde{Z}=\{\{\varphi_{1}(\mathbf{u}_{i})\}_{i=1}^{N}\mid\varphi_{1}\in\Phi_{1}\}$. We have
Proof [Proof of Lemma 13]
@@
Encoder and Decoder. The Encoder is case a), while the Decoder is case b) in Lemma 21.
@@
SSM. As discussed in Section 3.1, an SSM is equivalent to a linear transformation called its input-output map. Therefore, by Lemma 21, the SSM is $(\mu,0)$-RC in both cases, where $\mu$ is the operator norm of the input-output map. Combining this with Lemma 3 yields the result.
Remark 22
@@
As the value of $T$ is fixed, the input-output map can be described by the so-called Toeplitz matrix of the system. In this case, the operator norm equals to the appropriate induced matrix norm of the Toeplitz matrix. For the case of $T=\infty$, the input-output map still exists and is a linear operator. The proof of Lemma 21 holds in this case as well for operator norms.
Pooling. For any $Z\subseteq\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})$ we have
@@
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
Proof [Proof of Lemma 15] First of all, we show that the function $h:(\mathbb{R}^{2},\left\lVert\cdot\right\rVert_{2})\to(\mathbb{R},|\cdot|)$ defined as $h(\mathbf{x})=x_{1}\cdot\sigma(x_{2})$ is $\sqrt{2}(K+1)$-Lipschitz on a bounded domain, where $|x_{i}|\leq K$ for all $\mathbf{x}\in\mathbb{R}^{2}$ we consider. We will later specify the value of $K$ in relation to Assumption 9. By the sigmoid being 1-Lipschitz, we have
$\leq\sqrt{2}(K+1)\left\lVert\mathbf{x}-\mathbf{y}\right\rVert_{2}$
Second, we recall Corollary 4 in Maurer (2016).
Theorem 23 (Maurer (2016))
Let $\mathcal{X}$ be any set, $(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})\in\mathcal{X}^{N}$, let $\mathcal{F}$ be a set of functions $f:\mathcal{X}\to\ell^{2}(\mathbb{R}^{m})$ and let $h:\ell^{2}(\mathbb{R}^{m})\to\mathbb{R}$ be an $L$-Lipschitz function. Under $f_{k}$ denoting the $k$-th component function of $f$ and $\sigma_{ik}$ being a doubly indexed Rademacher variable, we have
We wish to apply Theorem 23 to GLU layers. For any $Z\subseteq\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})$ we have
$\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}GLU_{W}(\mathbf{z}_{i})\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
Now this is an alternative version of the Rademacher complexity, where we take the absolute value of the Rademacher average. In order to apply Theorem 23, we reduce the problem to the usual Rademacher complexity. In turn, we can apply the last chain of inequalities in the proof of Proposition 6.2 in Hajek and Raginsky (2019). Concretely, by denoting $\mathbf{O}=\{\mathbf{0}\}_{i=1}^{N}$ and noticing that $GLU_{W}(0)=0$, we have
Let $\mathbf{x}_{i}=i$, $i=1,\ldots,N$ and let =$\mathcal{F}=\{f_{W,\underline{z},k,j}\mid(W,\underline{z},k,j)\in\mathcal{W}\times(Z\cup\{0\})\times[T]\times[n_{u}]\}$ such that $f_{W,\underline{z},k,j}(\mathbf{x}_{i})=\begin{bmatrix}GELU(\mathbf{z}_{i}[k])^{(j)}&(W(GELU(\mathbf{z}_{i}[k])))^{(j)}\end{bmatrix}^{T}$ for $\underline{z}=\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z$. Since $Z\subseteq(B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(K_{\mathbf{u}}))^{N}$, it follows for all $\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z$ and for all $k\in\mathbb{N}$ that $\|\mathbf{z}_{i}[k]\|_{\infty}\leq K_{\mathbf{u}}$, and hence $|GELU(\mathbf{z}_{i}[k])^{(j)}|<K_{\mathbf{u}}$, leading to $|W(GELU(\mathbf{z}_{i}[k]))^{(j)}|<\sup\limits_{W\in\mathcal{W}}\|W\|_{\infty,\infty}\cdot K_{\mathbf{u}}$. In particular, $GLU^{(j)}_{W}(\mathbf{z}_{i})[k]=h(f_{W,\underline{z},k,j}(\mathbf{x}_{i}))=h|_{B}(f_{W,\underline{z},k,j}(\mathbf{x}_{i}))$, where $h|_{B}$ is the restriction of $h$ to $B=\{x\in\mathbb{R}^{2}\mid\|x\|_{\infty}<K\}$, and hence $h|_{B}$ is $\sqrt{2}(K+1)$-Lipschitz. Therefore we can set $K=\max\{K_{\mathbf{u}},\sup\limits_{W\in\mathcal{W}}\|W\|_{\infty,\infty}\cdot K_{\mathbf{u}}\}$.
We are ready to apply Theorem 23, together with the GLU definition and its $\sqrt{2}(K+1)$-Lipschitzness, we have
$=2\mathbb{E}_{\sigma}\left[\sup\limits_{f\in\mathcal{F}}\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}f(\mathbf{x}_{i})\right]\leq 4(K+1)\underbrace{\mathbb{E}_{\sigma}\left[\sup\limits_{f\in\mathcal{F}}\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}GELU(\mathbf{z}_{i}[k])^{(j)}\right]}_{A}$
$A=\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z\cup\{\mathbf{O}\}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}GELU(\mathbf{z}_{i})\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]=$
$\leq 4\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z\cup\{\mathbf{O}\}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]=4\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
and
$B=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in\{\mathbf{O}\}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W(GELU(\mathbf{z}_{i}))\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
$\leq 4\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z\{\mathbf{O}\}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
$=4\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
Here we used the linearity of $W$ and the exact same calculation as in the proof of Lemma 21.
By combining the inequalities above, it follows that
$16(K+1)\left(\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}+1\right)\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
Substituting the value of $K$ gives the result.
Proof [Proof of Lemma 16]
@@
MLP with sigmoid activations. Consider a single hidden layer $f(\mathbf{x})=\rho(g(\mathbf{x}))$, where $g(x)=W\mathbf{x}+\mathbf{b}$ is the preactivation and let $\mathcal{G}=\left\{g(x)=W\mathbf{x}+\mathbf{b}\mid W\in\mathcal{W},\mathbf{b}\in\mathcal{B}\right\}$ denote the set of possible preactivation functions. As compared to Definition 7, we omit the subscript from the notation of $g$. For an input sequence $\mathbf{z}\in\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})$ let $g(\mathbf{z})\in\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})$ mean that we apply $g$ for each timestep independently, i.e. $g(\mathbf{z})[k]=g(\mathbf{z}[k])$. We have
@@
$\mathbb{E}_{\sigma}\left[\sup\limits_{g\in\mathcal{G}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\rho(g(\mathbf{z}_{i}))\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
@@
$=\mathbb{E}_{\sigma}\left[\sup\limits_{(W,b)\in\mathcal{W}\times\mathcal{B}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}(W\mathbf{z}_{i}+\mathbf{b})\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})}\right]+\frac{1}{2\sqrt{N}},$
because the sigmoid is 1-Lipschitz and $\rho(0)=0.5$. Now we can apply Lemma 21 to get that
$\mathbb{E}_{\sigma}\left[\sup\limits_{(W,\mathbf{b})\in\mathcal{W}\times\mathcal{B}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}(W\mathbf{z}_{i}+\mathbf{b})\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})}\right]+\frac{1}{2\sqrt{N}}$
$\leq\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]+\frac{1}{\sqrt{N}}\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\mathbf{b}\right\rVert_{\infty}+\frac{1}{2\sqrt{N}}$
Therefore, the sigmoid MLP layer is $(K_{W},K_{\mathbf{b}}+0.5)$-RC. The restriction of an MLP to the ball $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ maps to the ball $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})}(1)$, because of the elementwise sigmoid activation.
@@
where we used that $\rho(x)=ReLU(x)$ is 1-Lipschitz and the same logic for the alternative definition of the Rademacher complexity as in the proof of Lemma 15, which results in a constant factor of 2. The constant 4 is then obtained by the additional constant factor 2 from Talagrand’s lemma. The rest of proof is identical to the sigmoid case.

### v2

@@ Appendix D Technical results on Rademacher contractions
$\leq\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\text{op}}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\ell^{p,p}_{T}(\mathbb{R}^{n_{u}})}\right]+\frac{1}{\sqrt{N}}\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\underline{\mathbf{b}}\right\rVert_{\ell^{q,q}_{T}(\mathbb{R}^{n_{v}})}$
where the first inequality follows from the triangle inequality, the first equality is the linearity of $W$, the second inequality follows from the definition of the operator norm, while the third and fourth inequalities refer only to the bias term and follow from the absolute homogenity of the norm and equation (3).
The values of $(p,q)$ influence the terms $\left\lVert W\right\rVert_{\text{op}}$ and $\left\lVert\underline{\mathbf{b}}\right\rVert_{\ell^{q,q}_{T}(\mathbb{R}^{n_{v}})}.$ As a result, for the separate cases of $a)$, $b)$ and $c)$ it is enough to separately bound these norms with constants similar to $K_{W}$. In the statement of the Lemma we assumed universal constants $K_{W}$ and $K_{\mathbf{b}}$ that do not depend on the values of $(p,q)$.
@@
We can see that the calculations hold if the transformations are restricted to the ball $B_{X_{1}}(r)$ for any choice of $X_{1}$ we consider. The radius can grow as
@@
$\left\lVert W(\mathbf{u})+\mathbf{b}\right\rVert_{\ell_{T}^{q,q}(\mathbb{R}^{n_{v}})}\leq\left\lVert W(\mathbf{u})\right\rVert_{\ell_{T}^{q,q}(\mathbb{R}^{n_{v}})}+\left\lVert\underline{\mathbf{b}}\right\rVert_{\ell_{T}^{q,q}(\mathbb{R}^{n_{v}})}\leq\left\lVert W\right\rVert_{\text{op}}\left\lVert\mathbf{u}\right\rVert_{\ell_{T}^{p,p}(\mathbb{R}^{n_{u}})}+\left\lVert\underline{\mathbf{b}}\right\rVert_{\ell_{T}^{q,q}(\mathbb{R}^{n_{v}})}.$
Remark D.2 is straightforward from the definitions of the considered norms.
D.1 Proofs for elements of deep SSMs
What we have left are the following. First, in light of the previous corollary, we need to show that each component of a deep SSM model is $(\mu,c)$-RC for some $\mu$ and $c$ w.r.t. compatible normed spaces. Second, we need to show that the Rademacher complexity of a $(\mu,c)$-RC model set are bounded in terms of $\mu$ and $c$. Lastly, there is a technical necessity due to the elementwise product in GLUs, namely we restrict the input space to the ball $B_{\ell^{2,2}_{T}(\mathbb{R}^{n_{in})}}(r)$, otherwise this restriction is not needed for SSMs without GLU activations. We start with the first one.
Lemma D.4.
Let $\mathcal{W}_{\text{Enc}}$, $\mathcal{W}_{\text{Dec}}$, and $\mathcal{E}$ denote some sets of parameters of some fixed Encoder, Decoder and SSM layers, respectively. Moreover, let $\{\mathcal{W}_{i}\times\mathcal{B}_{i}\}_{i=1}^{L+1}$ the parameter set of an MLP layer defined in Definition B.2, and let $\mathcal{W}_{\text{GLU}}$ be the parameter set of a GLU layer defined in Definition B.3. The corresponding function sets (in line with Assumption C.1) are
  • $\mathcal{F}_{\text{Enc}}=\{f^{\text{Enc}}=\langle W,\cdot\rangle\mid W\in\mathcal{W}_{\text{Enc}},\sup\limits_{W\in\mathcal{W}_{\text{Enc}}}\left\lVert W\right\rVert_{2,2}<K_{\text{Enc}}\}$,
  • $\mathcal{F}_{\text{Dec}}=\{f^{\text{Dec}}=\langle W,\cdot\rangle\mid W\in\mathcal{W}_{\text{Dec}},\sup\limits_{W\in\mathcal{W}_{\text{Dec}}}\left\lVert W\right\rVert_{\infty,\infty}<K_{\text{Dec}}\}$,
  • $\mathcal{F}_{\text{SSM}}=\{\mathcal{S}_{\Sigma}\mid\Sigma\in\mathcal{E},\sup\limits_{\Sigma\in\mathcal{E}}\left\lVert\Sigma\right\rVert_{p}<K_{p},p=1,2\}$,
  • $\mathcal{F}^{\rho}_{\text{MLP}}=\left\{f_{deep}=f_{W_{1},\mathbf{b}_{1}}\circ\ldots\circ f_{W_{L},\mathbf{b}_{L}}\circ g_{W_{L+1},\mathbf{b}_{L+1}}\middle|\begin{array}[]{l}(W_{i},\mathbf{b}_{i})\in\mathcal{W}_{i}\times\mathcal{B}_{i},\\ \sup\limits_{W\in\mathcal{W}_{i}}\left\lVert W\right\rVert_{\infty,\infty}<K_{W},\\ \sup\limits_{\mathbf{b}\in\mathcal{B}_{i}}\left\lVert\mathbf{b}\right\rVert_{\infty}<K_{\mathbf{b}},\\ 1\leq i\leq L+1\end{array}\right\}$,
  • $\mathcal{F}_{\text{GLU}}=\left\{f_{GLU}=GELU(\cdot)\odot\sigma\circ\langle W,GELU(\cdot)\rangle\middle|\begin{array}[]{l}W_{\text{GLU}}\in\mathcal{W}_{\text{GLU}},\\ \sup\limits_{W\in\mathcal{W}_{\text{GLU}}}\left\lVert W\right\rVert_{\infty,\infty}<K_{\text{GLU}}\end{array}\right\}$,
where $\mathcal{S}_{\Sigma}$ denotes the input-output map of the dynamical system $\Sigma$, $\rho$ is either the sigmoid or ReLU activations, and $\sigma$ denotes the sigmoid functions in the definition of GLU. Then all of these function sets are $(\mu,c)$-RC according to the following table, where for any Banach space $\mathcal{X}$, $B_{\mathcal{X}}(t)=\{x\in\mathcal{X}\mid\left\lVert x\right\rVert_{\mathcal{X}}\leq r\}$ denotes the ball of radius $r$ centered in zero for arbitrary $r$.
$\mu$ $c$ $X_{1}$ $X_{2}$
$\mathcal{F}_{\text{Enc}}$ $K_{\text{Enc}}$ 0 $B_{\ell_{T}^{2,2}(\mathbb{R}^{n_{\text{in}}})}(r)$ $B_{\ell_{T}^{2,2}(\mathbb{R}^{n_{u}})}(K_{\text{Enc}}r)$
$\mathcal{F}_{\text{Dec}}$ $K_{\text{Dec}}$ 0 $B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ $B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{\text{out}}})}(K_{\text{Dec}}r)$
$K_{\text{Dec}}$ 0 $B_{(\mathbb{R}^{n_{u}},\left\lVert\cdot\right\rVert_{\infty})}(r)$ $B_{(\mathbb{R}^{n_{\text{out}}},\left\lVert\cdot\right\rVert_{\beta})}(K_{\text{Dec}}r)$
$\mathcal{F}_{\text{SSM}}$ $K_{2}$ 0 $B_{\ell_{T}^{2,2}(\mathbb{R}^{n_{u}})}(r)$ $B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{y}})}(K_{2}r)$
$K_{1}$ 0 $B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ $B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{y}})}(K_{1}r)$
$\mathcal{F}_{\text{MLP}}^{ReLU}$ $4K_{W}(L+1)$ $4K_{\mathbf{b}}\sum\limits_{q=1}^{L}(4K_{W})^{q}$ $B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ $B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}\left(K_{W}^{L+1}r+K_{\mathbf{b}}\sum\limits_{q=1}^{L-1}K_{W}^{q}\right)$
$\mathcal{F}_{\text{MLP}}^{sigmoid}$ $K_{W}(L+1)$ $(K_{\mathbf{b}}+0.5)\sum\limits_{q=1}^{L}(K_{W})^{q}$ $B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ $B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}(K_{W}r+K_{\mathbf{b}})$
$\mathcal{F}_{\text{GLU}}$ $\begin{array}[]{l}16(r(K_{\text{GLU}}+1)^{2}\\ +K_{\text{GLU}}+1)\end{array}$ 0 $B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ $B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$
Furthermore, the operation of $f^{\text{Pool}}$, according the definition of deep SSMs, is $(1,0)$-RC between $X_{1}=B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ and $X_{2}=B_{(\mathbb{R}^{n_{u}},\left\lVert\cdot\right\rVert_{\infty})}(r)$.
Proof D.5.
@@ D.1 Proofs for elements of deep SSMs
Encoder and Decoder. The Encoder is case a), while the Decoder is case b) in Lemma D.1 along with Remark D.2.
@@
SSM. As discussed in Appendix B, an SSM is equivalent to a linear transformation called its input-output map. Therefore, by Lemma D.1, the SSM is $(\mu,0)$-RC in both cases, where $\mu$ is the operator norm of the input-output map. Combining this with Lemma A.3 yields the result.
Remark D.6.
@@
As the value of $T$ is fixed, the input-output map can be described by the so-called Toeplitz matrix of the system. In this case, the operator norm equals to the appropriate induced matrix norm of the Toeplitz matrix. For the case of $T=\infty$, the input-output map still exists and is a linear operator. The proof of Lemma D.1 holds in this case as well for operator norms.
Pooling. For any $Z\subseteq\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})$ we have
@@
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
@@
MLP with sigmoid activations. Consider a single hidden layer $f(\mathbf{x})=\rho(g(\mathbf{x}))$, where $g(x)=W\mathbf{x}+\mathbf{b}$ is the preactivation and let $\mathcal{G}=\left\{g(x)=W\mathbf{x}+\mathbf{b}\mid W\in\mathcal{W},\mathbf{b}\in\mathcal{B}\right\}$ denote the set of possible preactivation functions. As compared to Definition B.2, we omit the subscript from the notation of $g$. For an input sequence $\mathbf{z}\in\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})$ let $g(\mathbf{z})\in\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{v}})$ mean that we apply $g$ for each timestep independently, i.e. $g(\mathbf{z})[k]=g(\mathbf{z}[k])$. We have
@@
$\mathbb{E}_{\sigma}\left[\sup\limits_{g\in\mathcal{G}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\rho(g(\mathbf{z}_{i}))\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
@@
$=\mathbb{E}_{\sigma}\left[\sup\limits_{(W,b)\in\mathcal{W}\times\mathcal{B}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}(W\mathbf{z}_{i}+\mathbf{b})\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{v}})}\right]+\frac{1}{2\sqrt{N}},$
because the sigmoid is 1-Lipschitz and $\rho(0)=0.5$. Now we can apply Lemma D.1 (see Remark D.2) to get that
$\mathbb{E}_{\sigma}\left[\sup\limits_{(W,\mathbf{b})\in\mathcal{W}\times\mathcal{B}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}(W\mathbf{z}_{i}+\mathbf{b})\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{v}})}\right]+\frac{1}{2\sqrt{N}}$
$\leq\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]+\frac{1}{\sqrt{N}}\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\mathbf{b}\right\rVert_{\infty}+\frac{1}{2\sqrt{N}}$
Therefore, the sigmoid MLP layer is $(K_{W},K_{\mathbf{b}}+0.5)$-RC. The restriction of an MLP to the ball $B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ maps to the ball $B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{v}})}(1)$, because of the elementwise sigmoid activation. For the deep model the result is starightforward from Lemma 3.2 along with Lemma D.1, Remark D.2 and Remark E.1.
@@
where we used that $\rho(x)=ReLU(x)$ is 1-Lipschitz and the same logic for the alternative definition of the Rademacher complexity as in the proof of Lemma D.4, which results in a constant factor of 2. The constant 4 is then obtained by the additional constant factor 2 from Talagrand’s lemma. The rest of proof is identical to the sigmoid case.